OpenStudy (anonymous):

Let g(x) = integration from 0 to x of f(t) dt, where f(x) is the piecewise function below. f(x) = 0 x<0 x^2 0<(or equal to) x < 3 12-x 3<(or equal to) x < 10 2 x >(or equal to) 10 Given that f(x) is continuous, write g(x) as a piecewise function, in similar fashion to f(x).

OpenStudy (tkhunny):

One piece at a time: \(\int\limits_{-\infty}^0 0\; dt = 0\) \(\int\limits_0^3 t^{2}\;dt = What?\)

OpenStudy (anonymous):

they set it up with the same intervals like fromx<0 then 0<x<3 etc.

OpenStudy (anonymous):

so g(x)= ? for x<0 ....

OpenStudy (tkhunny):

g(x) = 0 for x < 0

OpenStudy (tkhunny):

Sorry, it should have been, for x < 0, \(g(x) = \int\limits_{-\infty}^{x}0\;dt\) Likewise, for 0 < x < 3, we have \(g(x) = \int\limits_{0}^{x}t^{2}\;dt\) With a little luck, it will be continuous at x = 0

OpenStudy (anonymous):

what about for intervals 3<(or equal to) x < 10 and x >(or equal to) 10 ? would it be more than integration?

OpenStudy (anonymous):

you have to add what you have accumulated so far

OpenStudy (anonymous):

\(g(x)=0\) if \(x<0\) \(g(x)=\frac{x^3}{3}\) if \(0\leq x<3\)

OpenStudy (anonymous):

since \(g(3)=\frac{3^3}{3}=9\) we have on \(3<x<10\) \[g(x)=12x-\frac{x^2}{2}+c\] and we find \(c\) by noting that \[g(3)=12\times 3-\frac{3^2}{2}+c=9\]

OpenStudy (anonymous):

thank you :)